OpenStudy (anonymous):

if a number is divisble by 7 but is not divisible by 21, then the number cannot be divisible by A. 2 B. 3 C. 5 D. 8 E. 10

OpenStudy (anonymous):

Think of it intuitively, What's 21/7?

OpenStudy (anonymous):

3 I guess but say I have 140 that isn't divisible by 3 or 8.

OpenStudy (anonymous):

But the question says CANNOT bc then I can contradict you and say 56 is divisible by 7 and not 21 but also divisible by 8 so 140 is just a case where it isn't

OpenStudy (anonymous):

It's 3 then right?

OpenStudy (anonymous):

Now look, since 21/7=3 Then really 21=7x3 Now if I have a number divisible by 7 and not 21 then it can't possibly be divisible by 3 Ex: 343/7=49 and this can be simplified to 7x7x7or 49x7 or 7x49 Now if we have 441 441/7=63 441=7x63=7x21x3=7x7x3x3=21x21

OpenStudy (anonymous):

Basically if you have a 7 in the list of factors and then a 3 in the list of factors of a number then the number is also divisible by 21cause then it's like 21x a number or 7x3 times a number, do you get it? And yes it's 3

OpenStudy (anonymous):

If the number was divisible by 21 then you could say that 7x3 but it cannot so you can't have a 3 in there or else you'll have 21 as factor

OpenStudy (anonymous):

Oh okay thank you so much :) I'm new to this thing so how do I honor you?

OpenStudy (anonymous):

If you click on my profile pic or my name you can go and press "view detailed profile" and whilst there you can leave a testimonial telling how good (or bad) of a tutor I was with you today

OpenStudy (anonymous):

OK I just clicked you as best response idk this is some confusing stuff with the medals.

OpenStudy (anonymous):

Ok, click on my picture or my name Then a box will appear, click view detailes profile, then on the left hand side click on testimonials and you'll see where it says to add one

OpenStudy (anonymous):

Sorry my bad, you have to click on my name and then click on become a fan in the bottom right and then the testimonial box will appear and you can type there

OpenStudy (anonymous):

Okayy! I got it!